OpenStudy (anonymous):

Another representation, in polar coordinates, of the point is (3,5pi/4) is (___,pi/4)

OpenStudy (anonymous):

since \(\frac{\pi}{4}\) is directly across the unit circle from \(\frac{\pi}{4}\) the coordinate \[(3,\frac{5\pi}{4})\] is the same as \[(-3,\frac{\pi}{4})\]

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

yw

OpenStudy (mathmale):

In other words, if you were to plot Satellite73's two points, they would be one and the same point.